At the beginning of 19X4, Beal Company adopted the following standards: Direct material (3 pounds &$2.50 per pound) 750 3750 Direct labor (5 hours $7.50 per hour) Factory overhead: Variable ($3.00 per dircet labor hour) Fixed ($4.00 per direct labor hour) 15.00 20.00 Standard cost per unit Normal volume per month is 40,000 direct labor hours Beals January 19X4 budget was based on normal volume. During January. Beal produced 7,800 units with records indicating the following: Direct material purchased Direct material used Direct labor Factory overhead 25.000 pounds e$2.60 23,100 pounds 40,100 hours @ $7.30 $300,000 Prepare a flexible budget for January 19X4 production costs based on actual production of 7,800 units For the month of January 19X4, compute the following variances, indicating whether each is favorable or unfavorable (a) (b) . 1. 2. Direct materials price variance, based on purchases Direct materials usage variance 3. Direct labor rate variance 4. Direct labor etficiency variance 5. Factory overhead spending variance 6. Variable factory overhead efficieney variance 7. Factory overhead volume variance Eastern Company manufactures special electrical equipment and parts Eastern uses standard costs with separate standards established for each product. A special transformer is manufac tured in the Transformer Department. Production volume is measured by direct labor hours in this department, and a flexible budget system is used to plan and control department overhead. Standard costs for the special transformer are determined annually in September for the coming
Answer #1

Part (a)

Flexible budget for the month ending 31st January 19X4

Please see the table. Please see the column titled "Working" for understanding the calculation of each item:

Parameter

Working

Value ($)

Direct materials

= 7800 units x 3 pounds x $ 2.5 / pound

58,500

Direct labor

= 7800 units x 5 hours x $ 7.5 / hour

292,500

Factory Overheads

- Variable

= 7800 units x 5 hours x $ 3 / hour

117,000

- Fixed

= 40000 hours x $ 4 / hour

160,000

Total budgeted production costs

Sum of all the above

628,000

Part (b)

Sub part: 1

Direct materials price variance based on materials purchased = (Actual cost per unit - Standard cost per unit) x actual material purchased = (2.60 - 2.50) x 25,000 = $ 2,500, Unfavorable (as actual cost per unit > standard cost per unit)

Sub part: 2

Direct materials usage variance = (Actual quantity used - standard quantity to be used) x standard cost per unit = (23,100 - 3 x 7,800) x 2.50 = - $ 750, Favorable (as actual quantity used is lower than the standard quantity)

Sub part: 3

Direct labor rate variance = (Actual labour rate - standard labour rate) x actual direct labor = (7.30 - 7.50) x 40,100 = - $ 8,020, Favorable (as actual labor rate < standard labor rate)

Sub part: 4

Direct labor efficiency variance = (Actual nos. of labour hours - standard nos. of labour hours) x standard labor rate = (40,100 - 5 x 7,800) x 7.50 = $ 8,250 Unfavorable (as actual nos. of labor hours > standard nos. of labor hours)

Sub part 5:

Factory overhead spending variance = Actual total factory overhead - budgeted total factory overhead at actual production volume

Actual total factory overhead = $ 300,000

Budgeted total factory overhead at actual production volume = Budgeted variable overhead at actual production volume + budgeted fixed overhead = 3 x 40,100 + 4 x 40,000 = $ 280,300

Factory overhead spending variance = Actual total factory overhead - budgeted total factory overhead at actual production volume = 300,000 - 280,300 = $ 19,700 Unfavorable (as actual overhead > budgeted overhead)

Sub part 6:

Variable factory overhead efficiency variance = Budgeted total factory overhead at actual hours - Budgeted total factory overhead at standard hours

Budgeted total factory overhead at actual hours = 280,300 (calculated in sub part 5 above)

Budgeted total factory overhead at standard hours = 117,000 + 160,000 = 277,000 (please see the production budget prepared in part (a)

Variable factory overhead efficiency variance = Budgeted total factory overhead at actual hours - Budgeted total factory overhead at standard hours = 280,300 - 277,000 = $ 3,300 Unfavorable

Sub part 7:

Factory overhead volume variance = Budgeted total factory overhead at standard hours - Applied total factory overhead

Budgeted total factory overhead at standard hours = 277,000 (as calculated in sub part 6 above)

Applied total factory overhead = actual production volume x standard labor hours x (Standard Variable overhead per direct labor hour + Standard fixed overhead per labor hour) = 7,800 x 5 x (3 + 4) = 273,000

Factory overhead volume variance = Budgeted total factory overhead at standard hours - Applied total factory overhead = 277,000 - 273,000 = 4,000 Unfavorable